OpenStudy (anonymous):

Solve: 2(x + 2) − (x + 1) = 8

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

2(x+2)-(x+1)=8 =>2x+4-x-1=8 =>x+3=8 x=5 yup :)